Consider a situation where a constant force of 25N acts on an object having a mass of 2kg for 3 seconds. What is the work done by the force? A) 150J

B) 50J C) 75J D) undetermined - not enough information E) none of the above
Answer #1

Impulse = Ft = 25 x 3 = 75 N*s

Momentum, p = mv. The 25 N force provides an acceleration of F / m (from F = ma) = 12.5 m/s^2
For 3 s so the increase in velocity, v is a 12.5 x 3 = 37.5 m/s. the momentum increases by a factor of 37.5.

w =mv^2/2 =1406.25 ans so option E is correct
Answer #2

FOR WORK DONE = F*S

HERE F IS GIVEN BUT for calculaton of S is not possible

as no initial velocity no accletration

so cant be determined